OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = -4, point (1, -3)

OpenStudy (anonymous):

since the line passes through (1,-3) it must be of the form (y+3)=m(x-1) or (y+3)=-4(x-1) or y=-4x+1